OpenStudy (anonymous):

point-slope equation for the line that passes through the point (5, -6) and is parallel to the line given by y = -10x + 3?

OpenStudy (anonymous):

point-slope equation for the line that passes through the point (5, -6) and is parallel to the line given by y = -10x + 3? A. y - 60 = -10(x - 5) B. y - 60 = -4x - 5 C. y - 6 = -10(x + 5) D. y + 6 = -10(x - 5)